# CamReasoner: Reinforcing Camera Movement Understanding via Structured Spatial Reasoning

### v2

@@ 4.5 Ablation Study
Table 3: Ablation of training stages. SFT enables structured reasoning, while RL further polishes performance.
Model SFT RL Binary Avg. VQA Avg.
Base (Qwen2.5-VL-7B) 35.8 60.3
CamReasoner-SFT 68.0 69.5
CamReasoner-7B 78.4 74.5
Table 4: Ablation study of the reward weight $\lambda$. A balanced weight of $\lambda=0.1$ achieves the optimal performance.
$\lambda$ Value 0.0 0.1 0.3 0.5
Binary Avg. 76.2 78.4 75.9 71.4
VQA Avg. 71.8 74.5 72.1 68.3
We evaluate the impact of training stages and the reward weight $\lambda$ on CameraBench. As shown in Table 3, the base model shows limited capability in interpreting complex cinematographic motions. Our SFT stage provides a foundational boost by aligning the model with the O-T-A reasoning paradigm, enabling it to decouple camera ego-motion from scene dynamics. The subsequent RL phase yields an additional 10.4% improvement in binary accuracy, demonstrating its effectiveness in optimizing decision-making and enforcing logical consistency across long-range temporal sequences.
Furthermore, Table 4 analyzes the sensitivity of $\lambda$, which balances accuracy ($r_{acc}$) and format ($r_{fmt}$) rewards. Performance peaks at $\lambda=0.1$, suggesting that while a small format reward is crucial for stabilizing the reasoning structure and ensuring parseable outputs, over-prioritizing formatting at higher values can detract from deep reasoning by constraining the model’s expressive capacity. This indicates that once the structural paradigm is established, the reinforcement signal should primarily focus on the correctness of the spatial-temporal deductions.
@@ 5 Conclusions
In this paper, we introduced CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task. By employing the O-T-A paradigm, the model explicitly decouples camera motion from complex backgrounds through spatial-temporal reasoning. Our construction of a Large-scale Inference Trajectory Suite, comprising 18k SFT and 38k RL samples, provides a foundational resource for teaching models cinematic logic. Experimental results show that CamReasoner-7B establishes a new state-of-the-art, achieving 78.4% accuracy in binary classification and outperforming GPT-4o by over 40%. Furthermore, our pioneering use of Reinforcement Learning for logical alignment effectively mitigates hallucinations in challenging scenarios like Confusable Motion. By bridging visual perception and structured deduction, CamReasoner offers a powerful foundation for future advancements in automated film analysis and controllable video synthesis.

### v3

@@ 4.5 Ablation Study
We conduct ablation studies on CameraBench to evaluate the contribution of each training stage and key RL hyperparameters.
Training Stages and RL Initialization Timing.
[…]
Reward Weight $\lambda$.
Table 4(b) (top) shows that removing the format reward ($\lambda$=0.0) causes a notable drop to 74.6%/70.2% due to occasional malformed outputs. Performance peaks at $\lambda$=0.1, while larger values progressively degrade accuracy as the model over-prioritizes formatting at the expense of reasoning depth.
GRPO Group Size $G$.
Table 4(b) (middle) shows that performance improves with larger groups, peaking at GG=8 (78.4%/74.5%). Further increasing to GG=16 yields no additional gains, suggesting eight samples already provide sufficiently stable advantage estimates.
Advantage Normalization.
Table 4(b) (bottom) validates EMA-GRPO over standard batch-level normalization (75.1%/71.5%). The optimal momentum $\alpha$=0.99 achieves the best performance by filtering high-frequency noise while remaining responsive to the evolving reward distribution.
5 Limitations
Several limitations of CamReasoner stem from objective resource and data constraints. First, due to computational budget, we validate CamReasoner only at the 7B scale; whether the O-T-A paradigm and RL-driven alignment yield comparable gains at larger (e.g., 72B) or smaller scales remains unexplored. Second, our training data derives entirely from the CameraBench distribution, which has limited coverage of certain motion categories—particularly confusable pairs such as dolly versus zoom—restricting the model’s exposure to diverse cinematic styles and domains; the 39.7% absolute accuracy on the out-of-domain ShotBench reflects this data scarcity. Third, the current pipeline lacks access to explicit geometric inputs such as monocular depth maps or camera intrinsics, which are not provided in existing video-QA benchmarks; integrating such signals could substantially improve disambiguation of visually degenerate motion pairs but requires either auxiliary annotation or additional foundation model inference at scale.
@@ 6 Conclusions
In this paper, we introduced CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task. By employing the O-T-A paradigm, the model explicitly decouples camera motion from complex backgrounds through spatio-temporal reasoning. Our construction of a Large-scale Inference Trajectory Suite, comprising 18k SFT and 38k RL samples, provides a foundational resource for instilling cinematic logic into MLLMs. Experimental results demonstrate that CamReasoner-7B, built upon Qwen2.5-VL-7B, substantially improves over its backbone, raising binary classification accuracy from 73.8% to 78.4% and VQA accuracy from 60.9% to 74.5%, consistently surpassing both proprietary models such as GPT-4o and strong open-source baselines. We hope CamReasoner can serve as a useful step toward bridging visual perception and structured deduction for camera movement understanding.
Appendix A Details of Curated Dataset
Figure 5: Distribution of camera movement categories in CamReasoning-SFT-18k. The dataset encompasses a diverse range of cinematographic motions, with a primary focus on dynamic rotations and stable static shots to foster robust reasoning capabilities.
To equip CamReasoner with a comprehensive understanding of cinematography, we curated the CamReasoning-SFT-18k dataset, which consists of 18,541 high-quality samples. As illustrated in Fig. 5, the dataset is categorized into four fundamental camera states to ensure balanced spatial-temporal learning:
  • Rotation (39.3%, 7284 samples): The largest portion of the dataset covers rotational movements, including Pan, Tilt, and Roll. This allows the model to learn the nuances of angular perspective shifts.
  • Static / Steadiness (27.5%, 5100 samples): A significant subset is dedicated to fixed-camera scenarios. These samples are crucial for teaching the model to distinguish between subject movement and actual camera ego-motion.
  • Zooming (18.3%, 3394 samples): This category focuses on focal length changes (Zoom In/Out), aiding the model in perceiving depth-related visual cues.
  • Track (14.9%, 2763 samples): This includes translational movements such as Truck, Pedestal, and Dolly, completing the model’s mastery of the basic cinematographic degrees of freedom.
This curated distribution ensures that the model does not overfit to a single motion type. By providing a substantial number of Static cases alongside Rotation and Track samples, we enforce the model’s ability to perform "deductive verification"—explicitly confirming the absence of movement when visual cues are ambiguous.
Appendix B Details of Evaluation Setting
Binary Classification
The original CameraBench evaluates binary classification using mean Average Precision (mAP) and ROC-AUC, which assess a model’s ability to rank samples by calibrated confidence scores. However, under our reasoning-based paradigm, models produce explicit reasoning chains followed by a final Yes/No verdict, where the resulting token-level probabilities tend to be poorly calibrated for score-based ranking. In contrast, accuracy directly measures whether a model reaches the correct final decision through its reasoning process, which better aligns with our evaluation goal. We therefore adopt an accuracy-based evaluation protocol, where all models are evaluated under the same setting for fair comparison, with discriminative models thresholded at 0.5.
VQA Evaluation Reformulation.
The original CameraBench VQA evaluation pairs each question with a positive and a negative video, and determines correctness by comparing the model’s yes/no confidence scores across the four possible (text, video) combinations. This protocol is designed to assess a model’s discriminative ability by measuring whether it assigns relatively higher scores to correct pairings, which is effective for evaluating confidence calibration across paired samples. However, in our reasoning-based setting, models produce explicit reasoning chains that analyze visual evidence before arriving at a final answer, making pairwise score comparison less naturally applicable. We therefore reformulate the task as multiple-choice: given a single video, the model must select which description correctly matches the observed camera movement from a set of candidates. This formulation complements the original by shifting the focus from relative score discrimination to explicit reasoning over visual content, requiring the model to articulate why one description applies over another. Both formulations evaluate camera movement understanding from different perspectives; our reformulation is specifically designed to better capture the reasoning process that our O-T-A paradigm emphasizes. The reformulated task is evaluated with standard accuracy under the same protocol for all models.
Appendix C Training Details
C.1 SFT Training
Table 5 indicates the main hyperparameters used during SFT training phase.
Table 5: Hyperparameter configuration for the SFT stage.
Hyperparameter Value
Model
Base model Qwen2.5-VL-7B-Instruct
Fine-tuning type Full
Vision tower Frozen
Multi-modal projector Frozen
Language model Trainable
Data
Max sequence length 16,384 tokens
Max training samples 20,000
Image max pixels 262,144
Video max pixels 16,384
Optimization
Optimizer AdamW
Learning rate $1\times 10^{-5}$
LR scheduler Cosine
Warmup ratio 0.03
Precision bf16
Training epochs 3
Infrastructure
Framework LLaMA-Factory
Parallelism DeepSpeed ZeRO-3
Per-device batch size 1
Gradient accumulation steps 2
Checkpoint interval 500 steps
C.2 GRPO Training
Table 6 indicates the main hyperparameters used during RL training phase.
Table 6: Full hyperparameter configuration for RL training.
Hyperparameter Value
Data
Max prompt length 16,384 tokens
Max response length 4,096 tokens
Video FPS 1.0
Max pixels 262,144
Optimization
Optimizer AdamW (bf16)
Learning rate $2\times 10^{-6}$
LR warmup ratio 0.0
Weight decay $1\times 10^{-2}$
Max gradient norm 1.0
RL Algorithm
Advantage estimator EMA-GRPO
KL penalty type low_var_kl
KL coefficient $1\times 10^{-2}$
Rollouts per prompt ($n$) 8
Rollout temperature 1.0
Rollout top-$p$ 1.0
Online filtering Top/bottom 1%
Filter key Accuracy
Reward weight $\lambda$ 0.1
Batch Sizes
Rollout batch size 128
Global batch size 32
Micro batch size (update) 1
Micro batch size (experience) 1
Validation batch size 128
Infrastructure
Precision bf16
FSDP Full shard
Gradient checkpointing Enabled
Padding free Enabled
Dynamic batching Enabled
Tensor parallelism 4
GPUs per node 4
Training epochs 2
Validation
Temperature 0.7
Top-pp 0.95
Samples per prompt 1
C.3 Computational Cost
[…]